OpenStudy (anonymous):

The sequence B(n) : n ≥ 0 is a geometric progression. If B(0) = 3 and B(1) = 6, complete the following: B(2)= B(3)= B(4)=

OpenStudy (anonymous):

You may assume that the sequence is of the form 3*(2^n) for n in N. So, B(2) = 3*(2^2) = 12. Can you figure out the rest? :-)

OpenStudy (anonymous):

B(3)=3*2^3=24......................IS IT CORRECT

OpenStudy (anonymous):

Yeah, that's the way to go :-)

OpenStudy (anonymous):

THANK U!

OpenStudy (anonymous):

No problem, mate :-)